OpenStudy (anonymous):

Find the roots of the polynomial equation. -x3 + 5x2 - 11x + 55 = 0

OpenStudy (rsadhvika):

factor first two terms, and last terms separately

OpenStudy (rsadhvika):

-x3 + 5x2 - 11x + 55 = 0 ^ ^

OpenStudy (anonymous):

Also, you can use the facts that the rational roots are the divisors of 55 so they are \[ \pm1, \pm 5, \pm 11 \] You find that the polynomial has a root x= 5, divide the polynomial by x-5 and you get the quotient then the answer. The method suggested by @rsadhvika is easier.

OpenStudy (anonymous):

Find the roots of the polynomial equation. -x3 + 5x2 - 11x + 55 = 0 A.\[i \sqrt{11}\] ,\[-i \sqrt{11}\] , –5 B.\[\sqrt{11}\] ,\[-\sqrt{11}\] , 5 C.\[i \sqrt{11}\] , –5 D.\[i \sqrt{11}\] , \[-i \sqrt{11}\], 5

OpenStudy (austinl):

\[-x^3 + 5x^2 - 11x + 55 = 0\] Perhaps this will help be a little more clear.

OpenStudy (anonymous):

im thinking that it is D

OpenStudy (rsadhvika):

why is that ? did u take time to read my or eliaasaab reply above ?

OpenStudy (whpalmer4):

You shouldn't have to think, you should be able to know :-)

OpenStudy (whpalmer4):

But yes, D is correct.

OpenStudy (whpalmer4):

\[-(i\sqrt{11})^3+5(i\sqrt{11})^2-11(i\sqrt{11})+55\]\[= -11i^3\sqrt{11}+55i^2-11i\sqrt{11}+55\]\[=-11(-1)i\sqrt{11}+55(-1)-11i\sqrt{11}+55\]\[=11i\sqrt{11}-55-11i\sqrt{11}+55 = 0\] conjugate root is similar \[-(5)^3+5(5)^2-11(5)+55 = -125+125-55+55=0\]